A certain bronze alloy containing 4% tin is to be added to 350kg of bronze containing 18% tin to produce a new bronze containing 15% tin. How many kilograms of the 4% bronze alloy are required?
Answer by stanbon(75887)

A certain bronze alloy containing 4% tin is to be added to 350kg of bronze containing 18% tin to produce a new bronze containing 15% tin. How many kilograms of the 4% bronze alloy are required?
----
Equation:
tin + tin = tin
0.04x + 0.18*350 = 0.14(x+350)
-----------------------------------
4x + 18*350 = 14x + 14*350
10x = 4*350
x = 4*35
x = 140 kg (amt of bronze alloy needed)
----------------
Cheers,
Stan H.
-----------